# Nucleon energy correlators as a probe of light-quark dipole operators at the EIC

### v1

@@ Abstract
Introduction  The search for new physics (NP) beyond the standard model (SM) has been a major focus of collider experiments such as the Large Hadron Collider (LHC). The lack of direct signals suggests that NP may emerge at an energy scale $\Lambda$ well beyond the current experimental reach. The standard model effective field theory (SMEFT) offers a systematic framework to describe such NP effects via higher-dimensional operators built from SM fields Buchmuller:1985jz ; Arzt:1994gp ; Grzadkowski:2010es . Among these, dimension-6 electroweak (EW) dipole operators involving light fermions are particularly intriguing for their chirality-flipping structure [^1] See, e.g., the quark dipole operator $(\bar{\mathcal{Q}}\sigma^{\mu\nu}d)HB_{\mu\nu}$ in Eq. (Nucleon energy correlators as a probe of light-quark dipole operators at the EIC). Alonso:2013hga ; Kley:2021yhn ; Aebischer:2021uvt ; Ardu:2025rqy ; Boughezal:2021tih ; Allwicher:2022gkm ; daSilvaAlmeida:2019cbr ; Li:2024iyj ; Gauld:2024glt ; Li:2025fom ; Escribano:1993xr ; Kopp:1994qv ; Boughezal:2023ooo ; Wen:2023xxc ; Wen:2024cfu ; Wen:2024nff , which is highly suppressed in the SM but essential for understanding electromagnetic dipole moments of leptons and baryons Hecht:2001ry ; Pitschmann:2014jxa ; Liu:2017olr ; Fukuyama:2012np ; Bhattacharya:2012bf ; Chupp:2017rkp ; Alarcon:2022ero ; Keshavarzi:2022kpc ; Gonzalez-Sprinberg:2024efj ; Muong-2:2021ojo ; Muong-2:2025xyk ; Aliberti:2025beg ; BESIII:2025vxm ; ACME:2018yjb ; Fan:2022eto ; Abel:2020pzs . However, constraints on these operators remain weak because the interference between dipole and SM amplitudes, scaling as ${\cal O}(\Lambda^{-2})$, typically vanishes in unpolarized observables Escribano:1993xr ; Kopp:1994qv ; daSilvaAlmeida:2019cbr ; Boughezal:2021tih ; Allwicher:2022gkm ; Li:2024iyj ; Gauld:2024glt ; Li:2025fom . Consequently, the leading effects arise at ${\cal O}(\Lambda^{-4})$, resulting in reduced sensitivity and contamination from dimension-8 operators.
@@
A new strategy to overcome this issue was recently proposed in Boughezal:2023ooo ; Wen:2023xxc , which leverages the transverse spin of fermions. Since transverse spin states are superpositions of helicity eigenstates, they enable interference between opposite-helicity components, restoring sensitivity to the dipole-SM interference at ${\cal O}(\Lambda^{-2})$. Building on this insight, several observables have been designed for the Electron-Ion Collider (EIC), the LHC, and future lepton colliders Wen:2023xxc ; Wen:2024cfu ; Wang:2024zns ; Boughezal:2023ooo ; Wen:2024nff . While transverse polarization of electron beams can be generated and controlled in accelerators, accessing transversely polarized quarks is more difficult due to color confinement. Existing approaches require non-perturbative correlations between quark spin and a transverse reference vector, either via nucleon polarization or via the transverse momenta of final-state hadrons, encoded in transversity parton distribution functions (PDFs) and dihadron fragmentation functions Boughezal:2023ooo ; Wen:2024cfu . Nevertheless, these methods either demand polarized nucleon beams, which compromise luminosity, or rely on semi-inclusive measurements that require particle identification and multi-particle tracking.
@@
To address these limitations, we propose nucleon energy correlators (NECs) as an innovative approach to probe light-quark EW dipole operators in inclusive deep inelastic scattering (DIS) with an unpolarized nucleon. Recently introduced in Liu:2022wop as an extension of energy correlators from final-state jets to nucleon structure, NECs enable the imaging of partonic dynamics through correlations between initial-state partons and energy flux in the target fragmentation region (TFR) Liu:2022wop ; Cao:2023oef ; Cao:2023qat ; Liu:2023aqb ; Liu:2024kqt ; Chen:2024bpj ; Mantysaari:2025mht ; Li:2023gkh ; Guo:2024jch ; Guo:2024vpe (see AbdulKhalek:2021gbh for a review). By exploiting these correlations, we construct a chiral-odd quark NEC that accesses the quark transverse spin inside an unpolarized nucleon. This NEC ensures a nonzero interference between the dipole and SM amplitudes, generating unique azimuthal asymmetries in the energy distribution in DIS. We demonstrate that these asymmetries exhibit clean and strong sensitivity to quark dipole couplings. Crucially, our method works without polarized nucleon beams and requires only inclusive calorimetric measurements, making it an exceptional observable for probing chirality-flipping interactions at both existing and future electron-ion colliders.
@@
Quark transversity NEC Let us begin by considering quark NECs in an unpolarized nucleon, which moves rapidly along the $\hat{z}$ direction with momentum $P$. The associated NECs are encoded in the following correlation function Liu:2022wop ; Chen:2024bpj :
@@
which describes the conditional probability of finding a quark with the longitudinal momentum fraction $x$ inside the target nucleon, given the observation of an energy flux at the solid angle $(\theta,\phi)$ in the TFR. $\theta$ and $\phi$ denote the polar and azimuthal angles, respectively. Formally, the quark NECs can be viewed as the collinear PDFs modified by the insertion of the energy flow operator ${\cal E}(\theta,\phi)$, which is defined through its action on hadronic states Sveshnikov:1995vi ; Bauer:2008dt
@@
where $E_{i}$ denotes the energy of hadron $i$, normalized by the nucleon energy $E_{N}$. Here, we use the light-cone coordinates defined as $a^{\mu}=(a^{+},a^{-},\vec{a}_{\perp})=\bigl{(}(a^{0}+a^{3})/{\sqrt{2}},(a^{0}-a^{3})/{\sqrt{2}},a^{1},a^{2}\bigr{)}$, and ${\cal L}_{n}$ denotes the light-cone gauge link.
@@
which is chiral-even and rotationally invariant in the transverse plane Liu:2022wop .
While prior studies have primarily focused on the chiral-even NECs Liu:2022wop ; Chen:2024bpj ; Liu:2024kqt ; Cao:2023oef ; Cao:2023qat ; Liu:2023aqb ; Mantysaari:2025mht (see a systematic classification in Chen:2024bpj ), we extend the formalism to the chiral-odd components. In particular, the quark transversity NEC can be obtained by using the projector $\Gamma^{\alpha}=i\sigma^{\alpha+}\gamma_{5}$ with $\alpha=1,2$, analogous to that used for transversity PDFs. However, unlike transversity PDFs, which requires a transversely polarized nucleon Jaffe:1991ra , NECs can access the quark transverse spin even in an unpolarized nucleon, owing to the azimuthal dependence of the observed energy flux.
To explicitly illustrate this feature, we introduce the vector $n_{T}^{\mu}=(0,0,\sin\theta\cos\phi,\sin\theta\sin\phi)$ to specify the transverse direction of the observed energy flux. Then the quark transversity NEC $h_{1}^{t,q}(x,\theta^{2})$ is defined as
$\frac{\epsilon_{\perp}^{\alpha\rho}n_{T,\rho}}{|n_{T}|}h_{1}^{t,q}(x,\theta^{2})={\cal M}^{[i\sigma^{\alpha+}\gamma_{5}]}(x,\theta,\phi)~,$ (4)
where $\varepsilon_{\perp}^{\mu\nu}=\varepsilon^{\mu\nu-+}$ with $\varepsilon^{0123}=1$. This equation shows that $h_{1}^{t}(x,\theta^{2})$ captures azimuthal correlations between the quark’s transverse polarization and the energy flow direction. Physically, it describes the difference in probability of finding a quark polarized along the transverse direction of the energy flow versus in the opposite direction. Being chiral-odd, the transversity NEC is off-diagonal in the helicity basis and naturally induces interference between amplitudes of opposite quark helicities. This provides a novel mechanism for generating chirality-flipping effects, enabling probes of dipole operators in unpolarized nucleon scatterings.
@@
Probing the quark dipole operators with transversity NEC To demonstrate how the transversity NEC probes the SMEFT dipole couplings, we consider the energy pattern within the TFR in inclusive DIS with an unpolarized nucleon, $l+p\rightarrow l^{\prime}+X$. Experimentally, this observable is determined from the angular distribution of the total energy deposited in the calorimeters. Following Refs. Liu:2022wop ; Chen:2024bpj , we define the energy pattern cross section as:
@@
We focus on the TFR, characterized by small polar angles ($\theta Q\ll Q$), where the energy flux originates from the fragmentation of target remnants. Following the general arguments in Cao:2023oef ; Chen:2024bpj , the energy pattern cross section in this region factorizes into hard partonic scattering coefficients and the non-perturbative NECs, defined by Eq. (1). Because NECs encapsulate all the information regarding the energy flow, including its azimuthal $\phi$ dependence, the $\phi$-independent cross section $\Sigma_{UU}$ factorizes with unpolarized NECs, such as $f_{1}$ given in Eq. (3). By contrast, the $\phi$-dependent terms $\Sigma^{\sin\phi}_{UU}$ and $\Sigma^{\cos\phi}_{UU}$ are expected to factorize with the transversity NEC $h_{1}^{t}$, which encodes azimuthal correlations as seen in Eq. (4). Remarkably, after factorization, the hard scatterings remain the same as in standard inclusive DIS process Chen:2024bpj .
Within the SM, the energy pattern is isotropic in the transverse plane. Since quark helicity flip in DIS hard scattering is suppressed by light quark masses, only the chiral-even unpolarized quark NEC $f_{1}(x,\theta^{2})$ contributes to $\Sigma(\theta,\phi)$ in the massless limit, while the chiral-odd contributions from the transversity NEC $h_{1}^{t}(x,\theta^{2})$ vanish to all orders in $\alpha_{s}$. The only non-zero contribution arises from the azimuthally independent cross section $\Sigma_{UU}$. At leading order in $\alpha_{s}$, the contribution from the photon exchanges takes the following form:
@@
Although the $\sin\phi$ and $\cos\phi$ azimuthal modulations are absent within the SM, they could arise from chirality-flipping NP effects, such as the following dimension-6 SMEFT dipole interactions Grzadkowski:2010es :
@@
We find that the $\sin\phi$-modulation is determined by the real parts of the quark dipole couplings $c_{q\gamma}$ and $c_{qZ}$ through the interference between photon and $Z$-boson exchanges. The corresponding cross section is given by:
@@
Our analysis can be readily generalized to include a longitudinally polarized electron beam, yielding the cross section $\Sigma_{LU}^{\sin\phi(\cos\phi)}$ factorized in terms of the transversity NEC $h_{1}^{t}$. Owing to the parity violation induced by electron polarization, the azimuthal modulations can be generated via $\gamma\gamma$ interference, without $Q^{2}/m_{Z}^{2}$ suppression present in the unpolarized case $\Sigma_{UU}^{\sin\phi(\cos\phi)}$. As indicated later, this leads to an enhanced sensitivity to the dipole couplings. Complete analytical results are provided in the Supplemental Material appendix .
@@
and similarly for $\mathscr{u}<0$. The polar-angle weight function $w(\theta)$ is introduced for later use. For a longitudinally polarized electron beam, one can analogously define the asymmetries $A_{LU}^{\mathscr{u}}$ by combining the calorimetric asymmetry with the longitudinal spin asymmetry, as we specify in appendix .
Numerical evaluation of these asymmetries requires non-perturbative input for the quark NECs $f_{1}$ and $h_{1}^{t}$. While the unpolarized quark NEC $f_{1}$ can be extracted directly from $\Sigma_{UU}$ in the TFR Liu:2022wop , one can also measure the transversity NEC $h_{1}^{t}$ without assuming SMEFT interactions by tagging an additional hadron in the current fragmentation region. Such measurements are feasible at facilities including HERA, JLab, and the EIC.
Although direct experimental determinations of NECs are not yet available, a remarkable correspondence has been established between $\theta$-weighted moments of the NECs and $k_{\perp}$-weighted moments of transverse-momentum-dependent parton distributions (TMDs), including the unpolarized quark NEC Liu:2024kqt . This correspondence can be straightforwardly extended to the transversity NEC. Specifically, we have
@@
where $\mathscr{f}_{1}$ and $\mathscr{h}_{1}^{\perp}$ are the unpolarized and Boer-Mulder quark TMDs, respectively. These relations allow us to infer the NECs from existing global fits of TMDs. In our analysis, we adopt the parameterization from in Barone:2009hw , which assumes a Gaussian form for the $k_{\perp}$-dependence.
To utilize the above relations, in Eq. (12) we choose the $\theta$-weight as $w(\theta)=|\sin\theta|$. The limits of the $\theta$-integration are determined by both theoretical and experimental considerations. The upper bound $\theta_{\text{max}}$ is imposed to ensure the validity of the factorization in the TFR Cao:2023oef ; Chen:2024bpj . This corresponds to a transverse momentum cutoff $k_{\perp}^{\text{max}}\sim Q\theta_{\text{max}}$ in the $k_{\perp}$-integrals in Eq. (13). Due to rapid falloff of the Gaussian TMDs at large $k_{\perp}$, this cutoff has negligible numerical impact on the integral. The lower bound $\theta_{\text{min}}$ is determined by the coverage of the calorimeter in the nucleon forward region. It imposes a minimal transverse-momentum threshold given by $k_{\perp}^{\text{min}}\approx E_{\text{min}}\theta_{\text{min}}$, where $E_{\text{min}}$ characterizes the energy resolution of calorimeters AbdulKhalek:2021gbh . Practically, we adopt $k_{\perp}^{\text{min}}\approx 0~\text{GeV}$ and $k_{\perp}^{\text{max}}=5~\text{GeV}$. With these inputs, we have evaluated numerically both $A_{UU}^{\sin\phi}$ and $A_{LU}^{\sin\phi}$ for the EIC kinematics, and the detailed results can be found in the Supplemental Material appendix .
Projected sensitivity at the EIC Although the EIC is primarily designed for exploring parton physics, its clean experimental environment and high luminosity also offer great potential for probing NP effects AbdulKhalek:2021gbh . To estimate the sensitivity to the dipole operators at the EIC, we perform a $\chi^{2}$ analysis based on the predicted asymmetries. We select a representative EIC configuration AbdulKhalek:2021gbh with a center-of-mass energy of $\sqrt{s}=105\ \text{GeV}$, which corresponds to the optimal integrated luminosity of $\mathcal{L}=100\,\text{fb}^{-1}$. The analysis is carried out in $Q$ bins for $Q\in[10,\,60]$ GeV and $x\in[0.01,\,0.5]$, incorporating an inelasticity cut $0.1\leq y\leq 0.9$. We assume the experimental value consistent with the leading-twist SM prediction, and is negligible. With this setup, the statistical uncertainty $\delta A_{i}\simeq\pi/(2\sqrt{\mathcal{E}_{\textrm{total}}})$ is dominant over the systematic uncertainties. Here, $\mathcal{E}_{\textrm{total}}=(\Sigma(\mathscr{u}>0)+\Sigma(\mathscr{u}<0))\times\mathcal{L}$ is the accumulated energy deposition.
@@
This correlation is significantly alleviated by incorporating the $A_{LU}^{\sin\phi}$ asymmetry. For the EIC, we assume $70\%$ polarization of the electron beam AbdulKhalek:2021gbh . Under this setup, single-operator constraints can reach the $\mathcal{O}(0.01)\sim\mathcal{O}(0.1)$ level. Importantly, while $A_{LU}^{\sin\phi}$ retains strong correlations between coefficients, the orientation of the corresponding blue band is different from that of $A_{UU}^{\sin\phi}$. Their combination thus greatly improves the parameter space resolution, confining the allowed region to a narrow area highlighted in red. Similar improvements are also observed in the constraints on $C_{dW}$-$C_{dB}$ and $c_{q\gamma}$-$c_{qZ}$ (see Supplemental Material appendix ). The constraints on the imaginary parts can be obtained analogously from the $\cos\phi$-asymmetry, yielding identical sensitivities. These results underscore the strong capability of the proposed asymmetries to unveil light-quark dipole interactions at the EIC.
@@
Figure 3: Comparison of projected $68\,\%$ C.L. constraints on $\text{Re}[C_{uB}]$ and $\text{Re}[C_{uW}]$ from HERA (green), the EIC (red), and the LHeC (blue), assuming $\Lambda=1$ TeV. The LHeC limits are scaled by a factor of 1010 for visibility. A detailed description of our analysis setup is provided in appendix .
@@
Besides the future EIC, the absence of a polarized nucleon beam requirement for these asymmetries also opens up opportunities to use available data from HERA. In Fig. 3, we compare the projected sensitivity from the combined-asymmetry analysis among the EIC, HERA Klein:2008di , and the LHeC LHeC:2020van . Although HERA benefits from higher-$Q^{2}$ coverage, its lower luminosity results in slightly weaker sensitivity compared to the EIC. However, it provides better discrimination between operators. With even higher energy and better luminosity, the LHeC could outperform both HERA and EIC by an order of magnitude.

### v2

# Nucleon Energy Correlators as a Probe of Light-Quark Dipole Operators at the Electron-Ion Collider

@@ Abstract
Introduction   The search for new physics (NP) beyond the standard model (SM) has been a major focus of collider experiments. The lack of direct signals suggests that NP may emerge at an energy scale $\Lambda$ well beyond the current experimental reach. The standard model effective field theory (SMEFT) offers a systematic framework to describe such NP effects via higher-dimensional operators built from SM fields Buchmuller and Wyler (1986); Arzt et al. (1995); Grzadkowski et al. (2010). Among these, dimension-6 electroweak (EW) dipole operators involving light fermions are particularly intriguing for their chirality-flipping structure [^1] See, e.g., the quark dipole operator $(\bar{\mathcal{Q}}\sigma^{\mu\nu}d)HB_{\mu\nu}$ in Eq. (Nucleon Energy Correlators as a Probe of Light-Quark Dipole Operators at the Electron-Ion Collider). Alonso et al. (2014); Kley et al. (2022); Aebischer et al. (2021); Ardu and Valori (2025); Boughezal et al. (2021); Allwicher et al. (2023); da Silva Almeida et al. (2019); Li et al. (2025b); Gauld et al. (2025); Li et al. (2025a); Escribano and Masso (1994); Kopp et al. (1995); Boughezal et al. (2023); Wen et al. (2023, 2024a, 2024b). Such effects are highly suppressed in the SM but essential for understanding anomalous magnetic moments (AMM) and $P$, $T$ violating electric dipole moments (EDM) of leptons and baryons  Hecht et al. (2001); Pitschmann et al. (2015); Liu et al. (2018); Fukuyama (2012); Bhattacharya et al. (2012); Chupp et al. (2019); Alarcon and others (2022); Keshavarzi et al. (2021); González-Sprinberg (2024); Abi and others (2021); Aguillard and others (2025); Aliberti and others (2025); Ablikim and others (2025); Andreev and others (2018); Fan et al. (2023); Abel and others (2020). However, constraints on these operators at colliders remain weak because the interference between dipole and SM amplitudes, scaling as ${\cal O}(\Lambda^{-2})$, typically vanishes in unpolarized observables Escribano and Masso (1994); Kopp et al. (1995); da Silva Almeida et al. (2019); Boughezal et al. (2021); Allwicher et al. (2023); Li et al. (2025b); Gauld et al. (2025); Li et al. (2025a). Consequently, the leading effects arise at ${\cal O}(\Lambda^{-4})$, resulting in reduced sensitivity and contamination from dimension-8 operators.
@@
A new strategy to overcome this issue was recently proposed in Boughezal et al. (2023); Wen et al. (2023), which leverages the transverse spin of fermions. Since transverse spin states are superpositions of helicity eigenstates, they allow interferences between opposite-helicity components, restoring sensitivity to the dipole-SM interference at ${\cal O}(\Lambda^{-2})$. Building on this insight, several observables have been designed for the Electron-Ion Collider (EIC) and future lepton colliders Wen et al. (2023, 2024a); Wang et al. (2024); Boughezal et al. (2023); Wen et al. (2024b). While transverse polarization of electron beams can be generated and controlled in accelerators, accessing transversely polarized quarks is more difficult due to color confinement. Existing approaches require non-perturbative correlations between quark spin and a transverse reference vector, either via nucleon polarization or via the transverse momenta of final-state hadrons, encoded in transversity parton distribution functions (PDFs) and dihadron fragmentation functions Boughezal et al. (2023); Wen et al. (2024a). Nevertheless, these methods either demand polarized nucleon beams, which compromise luminosity, or rely on semi-inclusive measurements that require particle identification and multi-particle tracking.
@@
To address these limitations, we propose nucleon energy correlators (NECs) as an innovative approach to probe light-quark EW dipole operators in inclusive deep inelastic scattering (DIS) with an unpolarized nucleon. Recently introduced in Liu and Zhu (2023) as an extension of energy correlators from final-state jets to nucleon structure, NECs enable the imaging of partonic dynamics through correlations between initial-state partons and final-state energy flux in the target fragmentation region (TFR) Liu and Zhu (2023); Cao et al. (2023, 2024); Liu et al. (2023); Liu and Zhu (2024); Chen et al. (2024); Mäntysaari et al. (2025); Li et al. (2023); Guo et al. (2025, 2024) (see Moult and Zhu (2025) for a review). By exploiting these correlations, we construct a chiral-odd quark NEC that accesses the quark transverse spin inside an unpolarized nucleon. This NEC ensures a nonzero interference between the dipole and SM amplitudes, generating unique azimuthal asymmetries in the energy distribution in DIS. We demonstrate that these asymmetries exhibit clean and strong sensitivity to the quark dipole couplings. Crucially, our method works without polarized nucleon beams and requires only inclusive calorimetric measurements, making it an exceptional observable for probing chirality-flipping interactions at both existing and future electron-ion colliders.
@@
Quark transversity NEC Let us begin by considering quark NECs in an unpolarized nucleon, which moves rapidly along the $\hat{z}$ direction with momentum $P$. The associated NECs are encoded in the following correlation function Liu and Zhu (2023); Chen et al. (2024):
@@
which describes the conditional probability of finding a quark with the longitudinal momentum fraction $x$ inside the target nucleon, given the observation of an energy flux at the solid angle $(\theta,\phi)$ in the TFR. $\theta$ and $\phi$ denote the polar and azimuthal angles, respectively. Formally, the quark NECs can be viewed as the collinear PDFs modified by the insertion of the energy flow operator ${\cal E}(\theta,\phi)$, which is defined through its action on hadronic states Sveshnikov and Tkachov (1996); Bauer et al. (2008)
@@
where $E_{i}$ denotes the energy of hadron $i$, normalized by the nucleon energy $E_{N}$. Here, we use the light-cone coordinates for a vector $a^{\mu}=(a^{+},a^{-},\vec{a}_{\perp})$ with $a^{\pm}=(a^{0}\pm a^{3})/\sqrt{2}$ , and ${\cal L}_{n}$ is the light-cone gauge link.
@@
which is chiral-even and rotationally invariant in the transverse plane Liu and Zhu (2023).
While prior studies have primarily focused on the chiral-even NECs Liu and Zhu (2023); Chen et al. (2024); Liu and Zhu (2024); Cao et al. (2023, 2024); Liu et al. (2023); Mäntysaari et al. (2025) (see a systematic classification in Chen et al. (2024)), we extend the formalism to the chiral-odd components. In particular, the quark transversity NEC can be obtained by using the projector $\Gamma^{\alpha}=i\sigma^{\alpha+}\gamma_{5}$ with $\alpha=1,2$, analogous to that used for transversity PDFs. The interpretation of quark transverse spin becomes explicit by rewriting $\Gamma^{\alpha}=\gamma^{+}({\cal Q}_{+}^{\alpha}-{\cal Q}_{-}^{\alpha})$, where ${\cal Q}_{\pm}^{\alpha}=\frac{1}{2}\big(1\mp\gamma^{5}\gamma^{\alpha}_{\perp}\big)$ projects onto quarks with transverse spin aligned or anti-aligned with the $\alpha_{\perp}$-direction Jaffe (1996). However, unlike transversity PDFs, which requires a transversely polarized nucleon Jaffe and Ji (1992), NECs can access the quark transverse spin even in an unpolarized nucleon, owing to the azimuthal dependence of the observed energy flux.
To explicitly illustrate this feature, we introduce a unit vector $\hat{n}_{T}^{\mu}=(0,0,\cos\phi,\sin\phi)$ to specify the transverse direction of the observed energy flux. By rotational covariance and parity property, the quark transversity NEC $h_{1}^{t,q}(x,\theta^{2})$ is parametrized as
$\epsilon_{\perp}^{\alpha\rho}{\hat{n}_{T,\rho}}h_{1}^{t,q}(x,\theta^{2})={\cal M}^{[i\sigma^{\alpha+}\gamma_{5}]}(x,\theta,\phi)~,$ (4)
where $\varepsilon_{\perp}^{\mu\nu}=\varepsilon^{\mu\nu-+}$ with $\varepsilon^{0123}=1$. This equation shows that $h_{1}^{t}(x,\theta^{2})$ captures azimuthal correlations between the quark’s transverse polarization and the energy flow direction. Physically, it describes the difference in probability of finding a quark polarized in the two opposite directions perpendicular to the transverse direction of the energy flow. Being chiral-odd, the transversity NEC is off-diagonal in the helicity basis and naturally induces interference between amplitudes of opposite quark helicities. This provides a novel mechanism for generating chirality-flipping effects, enabling probes of dipole operators in unpolarized nucleon scatterings.
@@
Probing the quark dipole operators with transversity NEC To demonstrate how the transversity NEC probes the SMEFT dipole couplings, we consider the energy pattern within the TFR in inclusive DIS with an unpolarized nucleon, $l+p\rightarrow l^{\prime}+X$. Experimentally, this observable is determined from the angular distribution of the total energy deposited in the calorimeters. Following Refs. Liu and Zhu (2023); Chen et al. (2024), we define the energy pattern cross section as:
@@
We focus on the TFR, characterized by small polar angles ($\theta P^{+}\ll Q$), where the energy flux originates from the fragmentation of target remnants. Following the general arguments in Cao et al. (2023); Chen et al. (2024), the energy pattern cross section in this region factorizes into hard partonic scattering coefficients and the non-perturbative NECs, defined by Eq. (1). Because NECs encapsulate all the information regarding the energy flow, including its azimuthal $\phi$ dependence, the $\phi$-independent cross section $\Sigma_{UU}$ factorizes with unpolarized NECs, such as $f_{1}$ given in Eq. (3). By contrast, the $\phi$-dependent terms $\Sigma^{\sin\phi}_{UU}$ and $\Sigma^{\cos\phi}_{UU}$ are expected to factorize with the transversity NEC $h_{1}^{t}$, which encodes azimuthal correlations as seen in Eq. (4). Remarkably, after factorization, the hard scatterings remain the same as in standard inclusive DIS process Liu and Zhu (2023); Cao et al. (2023); Chen et al. (2024).
Within the SM, the energy pattern is isotropic in the transverse plane. Since quark helicity flip in DIS hard scattering is suppressed by light quark masses, only the chiral-even unpolarized quark NEC $f_{1}(x,\theta^{2})$ contributes to $\Sigma(\theta,\phi)$ in the massless limit, while the chiral-odd contributions from the transversity NEC $h_{1}^{t}(x,\theta^{2})$ vanish to all orders in $\alpha_{s}$. The only non-zero contribution arises from the azimuthally independent cross section $\Sigma_{UU}$. At leading order in $\alpha_{s}$, the contribution from the photon exchanges takes the following form Liu and Zhu (2023):
@@
Although the $\sin\phi$ and $\cos\phi$ azimuthal modulations are absent within the SM, they could arise from chirality-flipping NP effects, such as the following dimension-6 SMEFT dipole interactions Grzadkowski et al. (2010):
@@
We find that the $\sin\phi$-modulation is linearly determined by the real parts of the quark dipole couplings $c_{q\gamma}$ and $c_{qZ}$ through the interference between photon and $Z$-boson exchanges. The corresponding cross section is given by:
@@
Our analysis can be readily generalized to include a longitudinally polarized electron beam, yielding the cross section $\Sigma_{LU}^{\sin\phi(\cos\phi)}$ factorized in terms of the transversity NEC $h_{1}^{t}$. Owing to the parity violation induced by electron polarization, the azimuthal modulations can be generated via $\gamma\gamma$ interference, without $Q^{2}/m_{Z}^{2}$ suppression present in the unpolarized case $\Sigma_{UU}^{\sin\phi(\cos\phi)}$. As indicated later, this leads to an enhanced sensitivity to the dipole couplings. Complete analytical results are provided in the Supplemental Material 1.
@@
and similarly for $\mathscr{u}<0$. The polar-angle weight function $w(\theta)$ is introduced for later use. For a longitudinally polarized electron beam, one can analogously define the asymmetries $A_{LU}^{\mathscr{u}}$ by combining the calorimetric asymmetry with the longitudinal spin asymmetry, as we specify in 1.
Numerical evaluation of these asymmetries requires non-perturbative input for the quark NECs $f_{1}$ and $h_{1}^{t}$. While the unpolarized quark NEC $f_{1}$ can be extracted directly from $\Sigma_{UU}$ in the TFR Liu and Zhu (2023), one can also measure the transversity NEC $h_{1}^{t}$ without assuming SMEFT interactions by tagging an additional hadron in the current fragmentation region. Such measurements are feasible at facilities including HERA, JLab, and the EIC.
Although direct experimental determinations of NECs are not yet available, a remarkable correspondence has been established between $\theta$-weighted moments of the NECs and $k_{\perp}$-weighted moments of transverse-momentum-dependent parton distributions (TMDs), including the unpolarized quark NEC Liu and Zhu (2024). This correspondence can be straightforwardly extended to the transversity NEC. Specifically, we have
@@
where $\mathscr{f}_{1}$ and $\mathscr{h}_{1}^{\perp}$ are the unpolarized and Boer-Mulder quark TMDs Boussarie and others (2023), respectively. These relations allow us to infer the NECs from existing global fits of TMDs. In our analysis, we adopt the parameterization from Barone et al. (2010), which assumes a Gaussian form for the $k_{\perp}$-dependence. The evolution effects are accounted for as described in 1.
To utilize the above relations, in Eq. (12) we choose the $\theta$-weight as $w(\theta)=|\sin\theta|$. The limits of the $\theta$-integration are determined by both theoretical and experimental considerations. The upper bound $\theta_{\text{max}}$ is imposed to ensure the validity of the factorization in the TFR Cao et al. (2023); Chen et al. (2024). This corresponds to a transverse momentum cutoff $k_{\perp}^{\text{max}}\sim Q\theta_{\text{max}}$ in the $k_{\perp}$-integrals in Eq. (13). Due to rapid falloff of the Gaussian TMDs at large $k_{\perp}$, this cutoff has negligible numerical impact on the integral. The lower bound $\theta_{\text{min}}$ is determined by the coverage of the calorimeter in the nucleon forward region. It imposes a minimal transverse-momentum threshold given by $k_{\perp}^{\text{min}}\approx E_{\text{min}}\theta_{\text{min}}$, where $E_{\text{min}}$ characterizes the energy resolution of calorimeters Abdul Khalek and others (2022). Practically, we adopt $k_{\perp}^{\text{min}}\approx 0~\text{GeV}$ and $k_{\perp}^{\text{max}}=5~\text{GeV}$. With these inputs, we have evaluated numerically both $A_{UU}^{\sin\phi}$ and $A_{LU}^{\sin\phi}$ for the EIC kinematics, and the detailed results can be found in the Supplemental Material 1.
@@
Projected sensitivity at the EIC To estimate the sensitivity to the dipole operators at the EIC, we perform a $\chi^{2}$ analysis based on the predicted asymmetries. We select a representative EIC configuration Abdul Khalek and others (2022) with a center-of-mass energy of $\sqrt{s}=105\ \text{GeV}$, which corresponds to the optimal integrated luminosity of $\mathcal{L}=100\,\text{fb}^{-1}$. The analysis is carried out in $Q$ bins for $Q\in[10,\,60]$ GeV and $x\in[0.01,\,0.5]$, incorporating an inelasticity cut $0.1\leq y\leq 0.9$. We assume the experimental value consistent with the leading-twist SM prediction, and is negligible. With this setup, the statistical uncertainty $\delta A_{i}\simeq\pi/(2\sqrt{\mathcal{E}_{\textrm{total}}})$ is dominant over the systematic uncertainties. Here, $\mathcal{E}_{\textrm{total}}=(\Sigma(\mathscr{u}>0)+\Sigma(\mathscr{u}<0))\times\mathcal{L}$ is the accumulated energy deposition.
@@
We observe that this correlation is significantly alleviated by incorporating the $A_{LU}^{\sin\phi}$ asymmetry. For the EIC, we assume $70\%$ polarization of the electron beam Abdul Khalek and others (2022). Under this setup, single-operator constraints can reach the $\mathcal{O}(0.01)\sim\mathcal{O}(0.1)$ level. Importantly, while $A_{LU}^{\sin\phi}$ retains strong correlations between coefficients, the orientation of the corresponding blue band is different from that of $A_{UU}^{\sin\phi}$. Their combination thus greatly improves the parameter space resolution, confining the allowed region to a narrow area highlighted in red. Similar improvements are also observed in the constraints on $C_{dW}$-$C_{dB}$ and $c_{q\gamma}$-$c_{qZ}$ (see Supplemental Material 1). The constraints on the imaginary parts can be obtained analogously from the $\cos\phi$-asymmetry, yielding identical sensitivities. The constraints are stable against different nonperturbative inputs of NEC, as we analyze in 1. These results underscore the strong capability of the proposed asymmetries to unveil light-quark dipole interactions at the EIC.
@@
Figure 3: Comparison of projected $68\,\%$ C.L. constraints on $\text{Re}[C_{uB}]$ and $\text{Re}[C_{uW}]$ from HERA (green), the EIC (red), and the LHeC (blue), assuming $\Lambda=1$ TeV. The LHeC limits are scaled by a factor of 1010 for visibility. A detailed description of our analysis setup is provided in 1.
@@
Besides the future EIC, the absence of a polarized nucleon beam requirement for these asymmetries also opens up opportunities to use available data from HERA. In Fig. 3, we compare the projected sensitivity from the combined-asymmetry analysis among the EIC, HERA Klein and Yoshida (2008), and the LHeC Agostini and others (2021). Although HERA benefits from higher-$Q^{2}$ coverage, its lower luminosity results in slightly weaker sensitivity compared to the EIC. However, it provides better discrimination between operators. With even higher energy and better luminosity, the LHeC could outperform both HERA and EIC by an order of magnitude.
Finally, let us comment on the photon-quark dipole coupling $c_{q\gamma}$ (given in Eq. (9)), which can be constrained by measuring hadronic magnetic and electric dipole moments at low energy. While high-precision neutron EDM measurements yield strong bounds on $\text{Im}[c_{q\gamma}]$Kley et al. (2022); Kumar and Mereghetti (2024), existing hadronic AMM measurements do not well constrain $\text{Re}[c_{q\gamma}]$, and thus $\text{Re}[C_{qW}]$ and $\text{Re}[C_{qB}]$. These real parts encode the $P$- and $T$-conserving NP effects and may produce sizable signatures without conflicting with neutron-EDM bounds. As shown in Fig. 3, our proposal provides sensitive limits on the real components; for explicit bounds on $\text{Re}[c_{q\gamma}]$, we refer the reader to the Supplemental Material 1. Consequently, this letter is dedicated to searching for these $P$- and $T$-conserving NP effects at $ep$ colliders.